# One-Shot Graph Representation Learning Using Hyperdimensional Computing

### v1

@@ Abstract
We present a novel, simple, fast, and efficient approach for semi-supervised learning on graphs. The proposed approach takes advantage of hyper-dimensional computing which encodes data samples using random projections into a high dimensional space (HD space for short). Specifically, we propose a Hyper-dimensional Graph Learning (HDGL) algorithm that leverages the injectivity property of the node representations of a family of graph neural networks. HDGL maps node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. Results of experiments with widely used benchmark data sets show that HDGL achieves predictive performance that is competitive with the state-of-the-art deep learning methods, without the need for computationally expensive training.
@@ 1 Introduction
Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations. Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction Li and Pi (2020); Kumar et al. (2020). Such problems are typically solved using graph embedding techniques Zhou et al. (2022); Cai et al. (2018); Xu (2021); Wang et al. (2022).
Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita (1997); Frasconi et al. (1998); Gori et al. (2005); Scarselli et al. (2009), which learn low-dimensional embeddings of nodes in a graph by aggregrating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning Kipf and Welling (2017); Veličković et al. (2018); Hamilton et al. (2017), because of their superior performance across a broad spectrum of applications Wu et al. (2020); Zhang et al. (2018); Zhou et al. (2022).
@@
Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact Desislavov et al. (2023); Strubell et al. (2020); de Vries (2023); Thompson et al. (2021). Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
@@
Hyperdimensional computing Kanerva (1988, 2009); Neubert et al. (2019); Thomas et al. (2021) offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector Kanerva (1988, 2009). All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al. (2019); Thomas et al. (2021). Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware Kleyko et al. (2023a, b). Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning Neubert et al. (2019); Kleyko et al. (2023b); Ge and Parhi (2020); Morris et al. (2022).
Against this background, we investigate Hyper-dimensional Graph Learner (HDGL), a one-shot approach to graph learning, that makes use of hyper-dimensional computing Kanerva (1988, 2009); Neubert et al. (2019); Thomas et al. (2021). Specifically, we outline a methodology for designing an HD mapping for node embedding that achieves competitive performance in semi-supervised node labeling tasks when compared to GNNs, which necessitate considerably more costly training. The key contributions of the paper are as follows:
  • We show that how to design a HD node embeddings that match the expressive power of GNN embeddings.
  • We propose HDGL, a graph learner that takes advantage of carefully designed hyperdimensional node embeddings to perform one-shot learning on graphs for semi-supervised node labeling tasks.
@@
The rest of the paper is organized as follows: Section 2 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN. Section 3 introduces HDGL, the proposed HD computing based one-shot algorithm for learning node embeddings. Section 4 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods. Section 5 concludes with a brief summary of the paper and an outline of some directions for further research.
@@ 2 Background and Preliminaries
2.1 Node Embedding Problem
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. Let $y_{v}\in\{1,\cdots,L\}$ be the the label associated with each $v\in\mathcal{V}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$. The goal is to learn a latent representation $\phi(v)$ of each node $v$ using the structure of $\mathcal{G}$ and $\mathbf{x}_{v}$ to accurately predict $y_{v}$.
@@ 2.2 Properties of HD Computing
Hyper-dimensional computing, originally introduced by Kanerva (1988) is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional static mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large Kanerva (1988, 2009). All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products Kanerva (1988); Neubert et al. (2019). The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations Kleyko et al. (2023a, b). In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$. We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ Kanerva (1988). This is also known as a blessing of HD spaces.
High Memory Capacity of HD Representation. As noted by Neubert et al. (2019), in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits Ahmad and Hawkins (2015); Neubert et al. (2019); Thomas et al. (2021). Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms Neubert et al. (2019).
@@ 2.3 HD Computing Operations
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler (1998) in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
Note that operations and vectors mentioned above have counterparts in the bipolar space $\{-1,1\}^{n}$.
@@ 2.4 GNN-based Node Embdeddings
Recall that GNN use the node features, together with the adjacency information to learn the embedding $\mathbf{h}_{v}$ of each node $v\in\mathcal{V}$. Most GNN use a neighborhood aggregation scheme where each node recursively aggregates the features of its neighbors to compute its feature vector Xu et al. (2019); Wu et al. (2020); Zhang et al. (2018); Zhou et al. (2022). After $k$ iterations of such aggregation, the node’s feature vector yields its embedding. More precisely, the latent representation of node $v$ at the $k$th iteration is given by:
$\mathbf{h}^{(k)}_{v}=\text{Agg}(\{\mathbf{W}^{(k)}\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\cup\{v\}\})$
@@ 2.5 Representational Power of GNN
Several authors have investigated the representational power of GNN from different perspectives, including their ability to distinguish between different graph structures Xu et al. (2019), e.g., using the Weisfeiler-Lehman (WL) test Weisfeiler and Leman (1968); Huang and Villar (2021) (which is known to decide graph isomorphism in general, barring a few exceptions, e.g., regular graphs), their ability to learn a set of functions on graphs Scarselli et al. (2009), and the equivalence of the preceding two perspectives Chen et al. (2019). The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors Weisfeiler and Leman (1968); Babai and Kucera (1979). The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors Huang and Villar (2021). Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into different representations. As shown by Xu et al. (2019); Morris et al. (2019), GNN are at most as powerful as the WL test in distinguishing graph structures; and for a GNN with a sufficient number of layers will be as powerful as the WL test if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme
$\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$
where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations Xu et al. (2019).

### v2

# Hyperdimensional Computing for Node Classification and Link Prediction

@@ Abstract
We introduce a novel method for transductive learning on graphs using hyperdimensional representations. The proposed approach encodes data samples using random projections into a very high-dimensional space (hyperdimensional or HD space for short). It obviates the need for expensive iterative training of the sort required by deep learning methods. Specifically, we propose a Hyperdimensional Graph Learning (HDGL) algorithm. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
We report results of experiments using widely used benchmark datasets which demonstrate that, on the node classification task, HDGL is competitive with the state-of-the-art GNN methods with respect to accuracy, at substantially reduced computational cost. Furthermore, HDGL is well-suited for class incremental learning where the model has to learn to effectively discriminate between a growing number of classes. Our experiments also show that the HD representation constructed by HDGL supports link prediction at accuracies comparable to that of DeepWalk and related methods, although it falls short of state-of-the-art Graph Neural Network (GNN) methods that rely on computationally expensive iterative training. We conclude that HDGL offers a computationally efficient alternative to graph neural networks for node classification, especially in settings that call for class-incremental learning or in applications that demand high accuracy models at significantly lower computational cost and learning time than possible with the state-of-the-art graph neural network methods.
@@ 1 Introduction
Node Classification and Link Prediction in Networks. Social networks, citation networks, etc., are naturally represented as graphs where nodes represent the entities that make up the network and edges between pairs of nodes encode relationships between entities. For example, in a social network, nodes represent individuals and links denote social connections; in a citation network, nodes represent articles, and (directed) links denote citations. Over the past decade, there has been considerable interest in graph learning problems such as node classification and link prediction [30, 29]. Such problems are typically solved using graph embedding techniques [63, 6, 60, 55].
1.1 Related Work
@@
Graph Neural Networks (GNNs), originally introduced by Sperduti and Starita [46], Frasconi et al. [13], Gori et al. [17], Scarselli et al. [44], which learn low-dimensional embeddings of nodes in a graph by aggregating information from local neighborhoods of the nodes, have recently emerged as the dominant approach to graph representation learning [26, 53, 18], because of their superior performance across a broad spectrum of applications [58, 62, 63]. Like other deep learning models, GNNs are data hungry, require extensive hyperparameter tuning, and achieve their superior performance at considerable computational cost, resulting in a large carbon footprint with attendant environmental impact [10, 47, 9, 50]. Hence, there is an urgent need for computationally more efficient, and consequently, environmentally more sustainable alternatives to deep learning models in general, with a specific emphasis on graph neural networks.
@@
Hyperdimensional computing [22, 23, 37, 49] offers a brain-inspired alternative to deep neural networks for artificial intelligence and machine learning applications, that maps each object using a sufficiently high-dimensional random encoding (HD encoding for short) to produce a binary or bipolar vector [22, 23]. All computations are performed on the objects thus encoded using simple operations, e.g., element-wise additions and dot products [22, 37, 49]. Computations on HD vectors, because they yield binary or bipolar vectors, can be realized using low-precision, fast, low-power, energy-efficient hardware [27, 28]. Hence, there is a growing interest in the use of HD computing in artificial intelligence and machine learning [37, 28, 15, 34].
Hyperdimensional encodings of graphs have been investigated by Poduval et al. [43] and Nunes et al. [38]. These works focus on performing graph operations, e.g., similarity based retrieval, graph classification, etc. In contrast, our focus is on methods that learn node-level representations for node classification and link prediction.
Hyperdimensional node embeddings have been explored by Kang et al. [24] and [31]. Specifically, Kang et al. [24] proposed RelHD, a novel processing-in-memory (PIM) hardware architecture [35] based on FeFET technology [4]; and Li et al. [31] introduced HyperNode for learning HD embeddings of nodes. However, both RelHD and HyperNode suffer from significant shortcomings associated with how they construct node representations from node features (see section 3.5 for details).
1.2 Overview and Contributions
We introduce a novel method for transductive learning on graphs using hyperdimensional representations. The proposed approach encodes data samples using random projections into a very high- dimensional space (hyperdimensional or HD space for short) [22, 23, 37, 49]. It obviates the need for expensive iterative training of the sort required by deep learning methods. The resulting latent node representations support both node classification and link prediction tasks, unlike typical deep learning methods, which often require separate models for these tasks.
Contributions. The key contributions of the paper are as follows:
  • We introduce HDGL, a method for learning HD node embeddings that match the expressive power of GNN embeddings. HDGL leverages the injectivity property of node representations of a family of Graph Neural Networks (GNNs) to map node features to the HD space and then uses HD operators such as bundling and binding to aggregate information from the local neighborhood of each node.
  • We show how the HD node embeddings learned by HDGL can be used to perform both semi-supervised node labeling and link prediction tasks. Thus, unlike deep learning methods usually require separate models for each of these tasks, HDGL produces a single model that can support both tasks.
@@ 1.2 Overview and Contributions
  • We present results experiments which show that HDGL offers a better alternative to graph neural networks for settings that call for class-incremental learning, where the model is required to effectively discriminate between a growing number of classes that are presented incrementally over time.
Organization. The rest of the paper is organized as follows: Section 3 reviews the HD computing framework, its basic operations, and their properties, and reviews the expressive power of GNN. Section 4 introduces HDGL, the proposed HD computing based algorithm for learning node embeddings. Section 5 describes results of extensive experiments that compare HDGL with the state-of-the-art GNN-based node embedding methods. Section 6 concludes with a brief summary of the paper and an outline of some directions for further research.
2 Problem Statement
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents set of undirected nodes, $\mathcal{E}$ represents the set of edges $\mathcal{E}\subseteq\{(u,v)|u,v,\in\mathcal{V}\}$. Let $N=|\mathcal{V}|$ denote the number of nodes in the graph. Supposed each node $v\in\mathcal{V}$ is described by a tuple of $d$ features $\mathbf{x}_{v}\in\mathbb{R}^{d}$. The $k$-hop neighbors of node $v$ in graph $\mathcal{G}$ are represented using multi-set $\mathcal{N}^{k}(v)$.
For the node label prediction problem in the transductive setting, let $y_{w}\in{1,\cdots,L}$ represent the label associated with each node $w\in\mathcal{W}$, where $\mathcal{W}\subset\mathcal{V}$. This implies that not all nodes in the graph are labeled. The goal is to learn a latent representation $\phi(v)$ for each node $v$, utilizing the structure of $\mathcal{G}$, the feature vector $\mathbf{x}_{v}$, and the labels $y_{w}$, to accurately predict the labels for a set of nodes that are unlabeled in the graph $\mathcal{G}$.In the transductive link prediction setting, the goal is to infer missing edges within the graph by leveraging the latent representations $\phi(v)$ of the nodes.
@@ 3.1 Properties of HD Computing
Hyper-dimensional computing, originally introduced by Kanerva [22] is a brain-inspired approach to cognitive computing that encodes each object $x\in\mathcal{X}$ using a high dimensional mapping $\phi:\mathcal{X}\rightarrow\mathcal{H}$, where $\mathcal{H}$ is typically $\{-1,1\}^{n}$ or $\{0,1\}^{n}$ where $n$ is sufficiently large [22, 23]. All computations are performed in $\mathcal{H}$, using simple operations, e.g., element-wise additions and dot products [22, 37]. The mapping $\phi$ is often random, which because its elements are low precision, lends itself to fast, low-power, energy-efficient hardware realizations [27, 28]. In this paper, we assume that =$\mathcal{H}=\{0,1\}^{n}$. We proceed to briefly summarize some of the key properties of HD representation and operations that can be used to compute with it.
Near Orthogonality of Random vectors in HD Space. HD encodings of objects are produced by sampling each dimension independently and uniformly from a distribution. Let $\mathbf{a}\in\{0,1\}^{n}$ and $\mathbf{b}\in\{0,1\}^{n}$ be two such HD vectors where $a_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$ and $b_{i}\sim Ber(p=0.5);\forall i\in\{1,\cdots,n\}$. We can then say that $\mathbf{a}$ and $\mathbf{b}$ are near orthogonal or $d_{H}[\mathbf{a},\mathbf{b}]$, the normalized hamming distance between $\mathbf{a}$ and $\mathbf{b}$, $\approx 0.5$ [22]. This is also known as a blessing of HD spaces.
High Memory Capacity of HD Representation. As noted by Neubert et al. [37], in the case of an $n$-dimensional binary HD space where each of the $n$-dimensional vectors is $f$-sparse, e.g., the number of nonzero entries is $fn$, the resulting HD space has capacity given by $\binom{n}{\lfloor f\cdot n\rfloor}$. For example, when $n=1000$, and $f=0.05$, the resulting HD space can store approximately $10^{80}$ or almost as many patterns as the number of atoms in the observable universe (which is estimated to be between $10^{78}$ and $10^{82}$).
Noise or Error Tolerance. Due to the distributed representation of data across numerous dimensions in HD spaces, errors in HD computing are limited to only a fraction of bits [1, 37, 49]. Thus, HD computing offers noise tolerance without the need for expensive error correction mechanisms [37].
@@ 3.2 HD Computing Operations
Rotation, also called permutation, denoted by $\Pi$, is a bitwise left shift operation $\Pi:\{0,1\}^{n}\rightarrow\{0,1\}^{n}$. A key property of Rotation is that rotation results in dissimilar/orthogonal vectors i.e $d_{H}[\mathbf{a},\Pi(\mathbf{a})]\approx 0.5$ and it is distributive over both Bundling and Binding and like binding, rotation is also reflective i.e $d_{H}[\mathbf{a},\mathbf{b}]\approx d_{H}[\Pi(\mathbf{a}),\Pi(\mathbf{b})]$. Rotation was introduced by Gayler [14] in order to protect or preserve information because information can be lost/cannot be distinguished because Binding is associative and commutative i.e:- $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})=(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$. In order to make the make a distinction between $(\mathbf{a}\otimes\mathbf{b})\otimes(\mathbf{c}\otimes\mathbf{d})$ and $(\mathbf{a}\otimes\mathbf{c})\otimes(\mathbf{b}\otimes\mathbf{d})$, rotation is used :- $(\mathbf{a}\otimes\mathbf{b})\otimes\Pi(\mathbf{c}\otimes\mathbf{d})$.
Note that the operations referenced earlier have equivalents in the bipolar space $\{-1,1\}^{n}$. Specifically, the Bundle operation can be represented as signed addition in the bipolar space, while the binding operation aligns with multiplication in the bipolar space.
@@ 3.4 Representational Power of GNN
Several authors have investigated the representational power of GNN, including their ability to distinguish between different graph structures [59], e.g., using the Weisfeiler-Lehman (WL) graph isomorphism test [56, 19] , or to learn a set of functions on graphs [44], or the equivalence of the preceding two notions of expressivity[8]. The WL test iteratively updates a node’s feature vector by aggregating the feature vectors of its network neighbors [56, 2]. The injectivity of the aggregation function used ensures that different neighborhoods of the nodes map to different feature vectors [19]. Alternatively, if the set of feature vectors of a node are represented as a multiset, the aggregation function must be able to map different multisets into distinct encodings. As shown by [59, 33], GNN are at most as powerful as the WL test in distinguishing graph structures if the mapping $\mathcal{A}:\mathcal{G}\rightarrow\mathbb{R}^{d}$ produced by the iterative aggregation scheme: $\mathbf{h}^{(k)}_{v}=g(\mathbf{h}^{(k-1)}_{v},f(\{\mathbf{h}^{(k-1)}_{j};j\in\mathcal{N}^{1}(v)\}))$ where $g(\cdot)$ and $f(\cdot)$ and $\mathcal{A}$’s graph-level readout are injective.
GNN generalize the WL test in that, in addition to discriminating different graph structures as they learn to map similar graph structures to similar representations [59].
3.5 Limitations of Previous HD Graph Methods
The HD Computing based graph-level learning methods introduced by Nunes et al. [38] and Poduval et al. [43] are designed for similarity-based graph retrieval and graph classification, as opposed to node classification or link prediction. Nunes et al. [38] constructs the Graph HD vector using node features derived using PageRank [5]. Poduval et al. [43] assign random HD-vectors to nodes and considers them as node features to build a Graph HD vector. Thus, both methods ignore node features present in the graph.
RelHD [24] constructs node representations for node classification tasks and uses the resulting representation to design a novel hardware chip design with a scalable processing in-memory (PIM) architecture based FeFET. However, RelHD has a significant limitation with respect how it encodes the node features into HD vectors. Specifically, it treats node features as dimension-wise indicator variables and computes the HD vector representation of each node by first generating $d$ random basis vector. Let $\Upsilon_{p}$ be such a random basis Hyper vector. Random basis vectors $\Upsilon_{p}$ are bundled/accumulated if feature at index $p$ is present i.e HD vector for a node $u$ is computed using basis vectors and raw features $\mathbf{x}_{u}=[x_{u1},\ldots,x_{ud}]$ using the following formula:- $\oplus_{p\in\{1,\ldots,d\}}(\mathbbm{1}(x_{up}\neq 0)\cdot\Upsilon_{p})$. Consequently, this method is not suitable for graphs with continuous or integer-valued node attributes in continuous form or even simple bag-of-words counts. Note that because HyperNode [31] also utilizes the same mechanism for constructing HD vector encodings of the nodes, it suffers from the same limitations as RelHD.
Furthermore, none of the aforementioned methods [38, 43, 24, 31] support link prediction on graphs.
In contrast, HDGL offers an effective and efficient method for constructing more expressive node embeddings that can accommodate not only binary or multi-valued, but also integer-valued or continuous node features. The resulting HD encodings of nodes support both node classification and link prediction.